OpenStudy (anonymous):

-How can x + 3 = -2x + 5 be set up as a system of equations? y = x + 13 y = -2x + 5 y = x + 3 y = -3x + 11 y = x + 12 y = -2x + 1 y = x + 3 y = -2x + 5

OpenStudy (anonymous):

make a guess, i bet you will be right!

OpenStudy (anonymous):

@satellite73 -.-

OpenStudy (anonymous):

c'mon, guess hint it is the only one with both \(x+3\) and \(-2x+5\)

OpenStudy (anonymous):

@satellite73 D

OpenStudy (anonymous):

got it on the first try!

OpenStudy (anonymous):

@satellite73 do you know this one Given the geometric sequence where a1 = 2 and the common ratio is 8, what is the domain for n? All integers All integers where n ≥ 1 All integers where n ≥ 0 All integers where n ≥ −1

OpenStudy (anonymous):

yes i do you want to guess again? you may be right hint: the first term is \(a_1\) look at the subscript

OpenStudy (anonymous):

\[\huge a_{\color{red}1}\]

OpenStudy (anonymous):

want to guess, or you want me to tell you?

OpenStudy (anonymous):

@satellite73 C

OpenStudy (anonymous):

no, it is B because you start at \(\color{red}1\) and \(n\geq 0\) starts at \(0\)

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

@satellite73 thank you

OpenStudy (anonymous):

yw got any others or is that all?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\(j(x)\) will have the highest, because if \(x=0\) then \(j(0)=10\times 2^0=10\times 1=10\) i.e the \(y\) intercept for \(j(x)\) is \(10\)

OpenStudy (anonymous):

\(f\) has \(6\) \(g\) has \(2\) and \(h\) has \(4\) so \(10\) is the largest

OpenStudy (anonymous):

ty:)@satellite73

OpenStudy (anonymous):

idk if you know this one The probability that a train leaves on time is 0.9. The probability that the train arrives on time and leaves on time is 0.36. What is the probability that the train arrives on time given that it leaves on time? 0.6 0.9 0.27 0.36

OpenStudy (anonymous):

@satellite73